# Discrete Approximate Circle Bundles

### v2

@@ 7.1 Optical Flow Patches
Many applications of computer vision and imaging involve data in the form of density or intensity functions on some region of space or time. The ambient dimension for such data is inevitably very large, and distances computed using the standard Euclidean metric often fail to reflect one’s intuition about how densities should compare (for example, two densities which differ by a small translation may be very far apart with respect to Euclidean distance). One would ideally opt for an Earth-mover-type distance such as the Wasserstein metric [RTG00], but these metrics are prohibitively expensive to compute, so an effective local-to-global approach to topological modeling which circumvents the need to construct a full distance matrix is highly desirable.
Consider the space $\mathcal{D}$ of real-valued functions on the unit ball $\mathbb{B}^{3}$. The space $\mathcal{D}$ has a natural right $SO(3)$-action defined by $(\rho\cdot A)(x)=\rho(A^{T}x)$ for all $\rho\in\mathcal{D}$, $A\in SO(3)$ and $x\in\mathbb{B}^{3}$. In particular, the orbit of any fixed $\rho_{0}\in\mathcal{D}$ is diffeomorphic to a quotient of $SO(3)$ by the corresponding stabilizer subgroup. For our demonstration, we sampled the $SO(3)$-orbit of a density $\rho_{\text{prism}}\in\mathcal{D}$ concentrated around the boundary of a triangular prism $P\subset\mathbb{B}^{3}$ (see Figure 10). Explicitly, $\rho_{\text{prism}}:\mathbb{B}^{3}\to\mathbb{R}$ is given by
@@
We generated a synthetic dataset by applying 25,000 random rotations to $\rho_{\text{prism}}$ and sampling the resulting densities on a lattice $L$ (in particular, we restricted a 32 x 32 x 32 lattice on $1,1]^{3}$ to $\mathbb{B}^{3}$). The sampled density values were then flattened to vectors $X_{\text{prism}}\subset\mathbb{R}^{32^{3}}$. Figure 11 shows the PCA cumulative explained variance $X_{\text{prism}}$, clearly illustrating that the embedding of the underlying 3-manifold is highly non-linear. Figure 12 shows a PCA visualization of a ’fat fiber’ of the dataset over a small neighborhood in $\mathbb{RP}^{2}$.
@@
To classify and coordinatize the resulting discrete approximate circle bundle, we constructed a cover $\mathcal{U}=\{U_{j}\}_{j=1}^{73}$ of $\mathbb{RP}^{2}$ and used the sparse circular coordinates algorithm to obtain a family of discrete approximate local trivializations, then used Construction 3.5 to obtain a witness $\Omega\in C^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ and compute characteristic class representatives. As suggested by Figure 12, we found that the standard Euclidean metric was sufficient to model the small portion of $X_{\text{prism}}$ in each $\pi^{-1}(U_{j})$.
Figure 13(a) shows the 1-skeleton of $\mathcal{N}(\mathcal{U})$. The node representing each $U_{j}\in\mathcal{U}$ is labeled by the rotated prism used to generate a sample $x_{j}\in X_{\text{prism}}$ whose base projection $\pi(x_{j})\in\mathbb{RP}^{2}$ was near the center of $U_{j}$. Note in particular that the prisms at nodes which are connected by an edge have similar axes of symmetry. The edge representing each 1-simplex $U_{jk}\in\mathcal{N}(\mathcal{U})$ is colored according to the value assigned by the Stiefel-Whitney class representative $\omega=\det_{*}(\Omega)\in Z^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ (blue for $1$, red for 1-1). We verified that $\omega$ was not a coboundary, confirming that one cannot consistently choose an orientation for the data (since $H^{1}(\mathbb{RP}^{2};\mathbb{Z}_{2})\cong\mathbb{Z}_{2}$, verifying non-triviality is sufficient to uniquely determine the orientation class). With respect to $\omega$, the twisted Euler number of $\Omega$ was found to be 3-3, in agreement with our theoretical model.
Using the algorithm described in Section 5, we computed the cobirth and codeath of each characteristic class representative (see Table 3). Each representative was found to be a cocycle on the entire nerve of $\mathcal{U}$, but the two cocycles only became coboundaries after 146 of the 216 1-simplices in $\mathcal{N}(\mathcal{U})$ were removed. One can interpret this result as a reflection of the compatibly non-orientable structures of the bundle and the base space. More specifically, the restriction of the underlying bundle $(\mathcal{O}_{\text{prism}},\pi,\mathbb{RP}^{2})$ to any non-nullhomotopic loop (e.g. an equator $\mathbb{RP}^{1}\subset\mathbb{RP}^{2}$) is non-orientable, hence isomorphic to the Klein bottle. Thus, the restriction of an orientation class representative $\omega\in Z^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ to any subcomplex of $\mathcal{N}(\mathcal{U})$ which includes a non-nullhomotopic loop must still be non-trivial.
@@
To illustrate, we used the coordinatization algorithm described in Section 6 to compute global coordinates for $X_{\text{prism}}$ compatible with the maximal stage $W\subset\mathcal{N}(\mathcal{U})$ of the weights filtration on which the characteristic class representatives became coboundaries. We isolated the set of densities $\widetilde{X}_{\text{prism}}\subset X_{\text{prism}}$ whose base point projections satisfied $|z|<0.1$, then projected these base points to the equator $\mathbb{RP}^{1}\subset\mathbb{RP}^{2}$. Figure 13(b) shows the 1-skeleton of $W$; the nodes which correspond to open sets containing the base projections of points in $\widetilde{X}_{\text{prism}}$ are highlighted. Figure 14 shows a sample of $\widetilde{X}_{\text{prism}}$ arranged according to assigned base angles in $\mathbb{RP}^{1}$ and assigned fiber angles in $\mathbb{S}^{1}$. Observe that the base angle assigned to each density roughly captures its axis of symmetry, and that this axis steadily traverses a rotation by $\pi$ as the base angle varies from 0 to $\pi$. The fiber angles assigned to densities in any narrow vertical strip roughly parametrize a rotation by $\frac{2\pi}{3}$ about a common axis of symmetry. On the other hand, careful inspection shows that the direction of rotation abruptly flips near the dotted vertical line, reflecting a location in the nerve where a cut was made before coordinatization. This also confirms the Klein bottle structure of the restricted bundle underlying $\widetilde{X}_{\text{prism}}$.
@@
Figure 14: A collection of coordinatized samples from $\widetilde{X}_{\text{prism}}$ arranged according to base projection angle (x) and assigned fiber coordinate (y). The dotted vertical line indicates an orientation flip in the coordinate system.
Finally, we used the pipeline described in Section 6 to construct a point cloud embedding $F:X_{\text{prism}}\to V(2)$ which respects the global topology implied by the characteristic classes (recall these were well-defined on the entire nerve $\mathcal{N}(\mathcal{U})$). In this sense, our pipeline serves a similar function as Isomap, but with the preservation of topological (rather than geometric) features in mind. The mean projection errors for different projection dimensions are shown in Figure 15 .
@@
7.3 Geometric Signal Unwrapping
Our final implementation provides an example of how the techniques and algorithms developed here can be integrated with existing topological modeling tools to form a larger local-to-global TDA framework. We consider a second dataset of discretized densities, sampled from the $SO(3)$-orbit of the density $\rho_{\text{star}}\in\mathcal{D}$ shown in Figure 16. Note that $\rho_{\text{star}}$ has a single 5-fold rotational symmetry in the $yz$-plane, so $\mathcal{O}_{\text{star}}$ is homeomorphic to the lens space $SO(3)/\mathbb{Z}_{5}\cong L(10,1)$. Applying the same projection map as we used for $\mathcal{O}_{\text{prism}}$ yields a bundle $(\mathcal{O}_{\text{star}},\pi,\mathbb{RP}^{2})$ with fiber $\mathbb{S}^{1}\coprod\mathbb{S}^{1}$.
Figure 16: A sample of 3D densities from $X_{\text{star}}$. Each density is represented by its 2D projection along the $x$-axis as well as the corresponding rotated triangle mesh used to generate the density.
Figure 17: A PCA visualization of the data in a ’fat fiber’ of $X_{\text{star}}$ (right) over a small neighborhood in $\mathbb{RP}^{2}$ (left), portrayed here by a hemisphere with boundary identifications.
[…]
As in the previous experiment, we sampled 25,000 densities from $\mathcal{O}_{\text{star}}$ and formed a dataset $X_{\text{star}}\subset\mathbb{R}^{32^{3}}$ from the discretizations. We used a Mapper-like pipeline to cluster the data $X_{j}=X_{j}^{+}\coprod X_{j}^{-}$ over each open set $U_{j}$ in the same cover of $\mathbb{RP}^{2}$ (the sign labels for each pair of local clusters $X_{j}^{\pm}$ were assigned arbitrarily). Figure 18(a) shows a graph $G$ whose nodes are the local clusters $\{X_{j}^{\pm}\}_{j=1}^{73}$ and whose edges represent intersections between clusters. Edges between local clusters with the same sign label are shown in blue, and intersections between local clusters with opposite sign labels are shown in red. Figure 18(b) shows the 1-skeleton of $\mathcal{N}(\mathcal{U})$ with edges colored according to the values assigned by the induced connectivity cocycle $\nu\in Z^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ (blue for $1$ and red for 1-1). Note that each edge in Figure 18(b) corresponds to a pair of edges in Figure 18(a) of the same color. Figure 19 gives a more detailed visualization of $G$; each local cluster $X_{j}^{\pm}$ is labeled by the rotated star pyramid used to generate a sample density it contains. For readability, edges connecting nodes with opposite sign labels are not shown explicitly, but their presence is indicated by light gray lines. Lastly, observe that $G$ can be canonically identified with the 1-skeleton of the nerve of a good open cover $\widetilde{\mathcal{U}}$ of $\mathbb{S}^{2}$.
We verified that $\nu$ was a non-trivial cocycle, confirming that $X_{\text{star}}$ could not be globally separated into two connected components and that $\pi|_{X_{\text{star}}}$ could instead be lifted to a map $\tilde{\pi}:X_{\text{star}}\to\mathbb{S}^{2}$. To explicitly construct $\tilde{\pi}$, we selected a random $v_{\rho}\in X_{1}^{+}$ and chose an arbitrary lift $\tilde{\pi}(v_{\rho})\in\mathbb{S}^{2}$ of $\pi(v_{\rho})\in\mathbb{RP}^{2}$. For every other $v_{\rho^{\prime}}\in X_{1}^{+}$, we then defined $\tilde{\pi}(v_{\rho^{\prime}})$ to be the lift‘ of $\pi(v_{\rho^{\prime}})$ whose inner product with $\tilde{\pi}(v_{\rho})$ is positive. Similarly, for each $v_{\rho^{\prime}}\in X_{1}^{-}$, we defined $\tilde{\pi}(v_{\rho^{\prime}})$ so that $\tilde{\pi}(v_{\rho^{\prime}})\cdot\tilde{\pi}(v_{\rho})<0$. We then propagated this choice of lift across the entire dataset using a breadth-first search of $G$. Finally, we used the sparse circular coordinates algorithm to obtain a family of discrete approximate local trivializations for $X_{\text{star}}$ over $\mathbb{S}^{2}$, then used the algorithms described in Section 4 to obtain representatives of the orientation and twisted Euler classes. The results are summarized in Table 2. Note the Stiefel-Whitney class representative we obtained was trivial, and we obtained a twisted Euler number of $10$, as expected.
(a) Local Cluster Graph $G$
(b) Connectivity Cocycle $\nu$
Figure 18: Visualizations of (a) the graph $G$ which represents the connectivity relations among local clusters in $X_{\text{star}}$. (b) The 1-skeleton of the nerve $\mathcal{N}(\mathcal{U})$ of the open cover of $\mathbb{RP}^{2}$ used to analyze $X_{\text{star}}$ (right). Each 1-simplex of $\mathcal{N}(\mathcal{U})$ is colored according to the assigned value of the connectivity class representative $\nu$ (blue for 1, red for -1), and the pair of corresponding edges in $G$ are assigned this same color.
Figure 19: A detailed visualization of the local cluster graph $G$ which double covers the 1-skeleton of $\mathcal{N}(\mathcal{U})$. Nodes corresponding to local clusters labeled with a ++ are shown on the left, and the corresponding clusters labeled by a - are shown on the right. Edges which connect local clusters with the same sign label are shown in blue. For readability, edges which connect clusters with opposite sign labels are not shown explicitly, but are indicated with gray lines. Note the diagram above can equivalently be interpreted as representing the nerve of the cover $\widetilde{\mathcal{U}}$ of $\mathbb{S}^{2}$ which we ultimately used to model $X_{\text{star}}$ as a discrete approximate circle bundle.
The process of lifting a bundle with fiber $\mathbb{S}^{1}\coprod\mathbb{S}^{1}$ to a genuine circle bundle over a suitable covering space can be viewed as a form of geometric signal unwrapping. In other fields, classical signal unwrapping techniques (e.g., in interferometry, MRI, or phase retrieval) are used to resolve ambiguities in angular measurements (see, for example, [GP98]); our construction provides a topological analog in which the discrete fiber ambiguity is resolved via a covering, yielding a continuous, globally well-defined structure.
@@ 8 Conclusion and Future Work
A natural direction for future research is to establish a recovery theorem providing conditions under which a finite metric space $X$ and a feature map $\pi:X\to B$ uniquely determine a true circle bundle, independent of the choice of discrete, approximate local trivializations.
@@ A Technical Details
A.1 The Group $\text{Homeo}(\mathbb{S}^{1})$
In this section, we prove some technical results about $\text{Homeo}(\mathbb{S}^{1})$. We denote the connected components of orientation-preserving and -reversing homeomorphisms by $\text{Homeo}_{+}(\mathbb{S}^{1})$ and $\text{Homeo}_{-}(\mathbb{S}^{1})$, respectively. A model for the universal cover $\widetilde{\text{Homeo}}_{\pm}(\mathbb{S}^{1})$ of each component is given by
$\widetilde{\text{Homeo}_{\pm}(\mathbb{S}^{1})}=\{f\in\text{Homeo}(\mathbb{R}):f(x+1)=f(x)\pm 1\}$ (29)
We then define $\widetilde{\text{Homeo}(\mathbb{S}^{1})}=\widetilde{\text{Homeo}_{+}(\mathbb{S}^{1})}\coprod\widetilde{\text{Homeo}_{-}(\mathbb{S}^{1})}$.
We denote the group of isometries of $\mathbb{S}^{1}$ by $\text{Iso}(\mathbb{S}^{1})<\text{Homeo}(\mathbb{S}^{1})$ and denote the connected components by $\text{Iso}_{\pm}(\mathbb{S}^{1})$. We have a natural identification of $\text{Iso}_{+}(\mathbb{S}^{1})$ with $\mathbb{S}^{1}$ itself via the map $R^{+}:\mathbb{S}^{1}\to\text{Iso}_{+}(\mathbb{S}^{1})$ which sends $e^{i\theta}$ the counter-clockwise rotation $R^{+}_{\theta}$ (equivalently, $R^{+}_{\theta}$ denotes left-translation in $\mathbb{S}^{1}$ by $e^{i\theta}$). Similarly, every orientation-reversing isometry is expressible as $R_{\theta}^{-}$ for some $\theta$, where $R_{\theta}^{-}(z)=e^{i\theta}\bar{z}$ for all $z\in\mathbb{S}^{1}$. We can model the universal covers $\widetilde{\text{Iso}}_{\pm}(\mathbb{S}^{1})$ as subsets of $\widetilde{\text{Homeo}}_{\pm}(\mathbb{S}^{1})$:
$\widetilde{\text{Iso}}_{\pm}(\mathbb{S}^{1})=\{T^{\pm}_{a}(x)=\pm x+a\ :\ a\in\mathbb{R}\}$ (30)

### v3

@@ 7.2 Folded Klein Bottle
Many applications of computer vision and imaging involve data in the form of density or intensity functions on some region of space or time. The ambient dimension for such data is inevitably very large, and distances computed using the standard Euclidean metric often fail to reflect one’s intuition about how densities should compare (for example, two densities which differ by a small translation may be very far apart with respect to Euclidean distance). One would ideally opt for an Earth-mover-type distance such as the Wasserstein metric [22], but these metrics are prohibitively expensive to compute, so an effective local-to-global approach to topological modeling which circumvents the need to construct a full distance matrix is highly desirable. Furthermore, for our experiment below, we sample from a 3-manifold embedded in a space of 3D densities, and persistence diagrams for dimension 3 are computationally intractable using any metric; the only way to certify the global topology is with local computations.
Consider the space $\mathcal{D}$ of real-valued functions on the unit ball $\mathbb{B}^{3}$. The space $\mathcal{D}$ has a natural right $SO(3)$-action defined by $(\rho\cdot A)(x)=\rho(A^{T}x)$ for all $\rho\in\mathcal{D}$, $A\in SO(3)$ and $x\in\mathbb{B}^{3}$. In particular, the orbit of any $\rho_{0}\in\mathcal{D}$ is diffeomorphic to a quotient of $SO(3)$ by the corresponding stabilizer subgroup. For our demonstration, we sampled the $SO(3)$-orbit of a density $\rho_{\text{prism}}\in\mathcal{D}$ concentrated around the boundary of a triangular prism $P\subset\mathbb{B}^{3}$ (see Figure 14). Explicitly, $\rho_{\text{prism}}:\mathbb{B}^{3}\to\mathbb{R}$ is given by
@@
We generated a synthetic dataset by applying 5,000 random rotations to $\rho_{\text{prism}}$ and sampling the resulting densities on a lattice $L$ (in particular, we restricted a 32 x 32 x 32 lattice on $1,1]^{3}$ to $\mathbb{B}^{3}$). The sampled density values were then flattened to vectors $X_{\text{prism}}\subset\mathbb{R}^{32^{3}}$. Figure 15 shows the PCA cumulative explained variance $X_{\text{prism}}$, clearly illustrating that the embedding of the underlying 3-manifold is highly non-linear. Figure 16 shows a PCA visualization of a ’fat fiber’ of the dataset over a small neighborhood in $\mathbb{RP}^{2}$.
@@
To classify the resulting discrete approximate circle bundle and infer the topology of the 3-manifold underlying $X_{\text{prism}}$, we constructed a cover $\mathcal{U}=\{U_{j}\}_{j=1}^{60}$ of $\mathbb{RP}^{2}$; as suggested by Figure 16, we found (to our surprise) that PCA was sufficient to compute local circular coordinates on the small portion of $X_{\text{prism}}$ in each $\pi^{-1}(U_{j})$.
Figure 17(a)(a) shows the 1-skeleton of $\mathcal{N}(\mathcal{U})$. Each vertex $j\in\mathcal{N(U)}$ is labeled by the rotated prism used to generate a sample $x_{j}\in X_{\text{prism}}$ whose base projection $\pi(x_{j})\in\mathbb{RP}^{2}$ was near the center of $U_{j}$. Note in particular that the prisms at nodes which are connected by an edge have similar axes of symmetry. The edge representing each 1-simplex $(jk)\in\mathcal{N}(\mathcal{U})$ is colored according to the value assigned by the Stiefel-Whitney class representative $\omega=\det_{*}(\Omega)\in Z^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ (blue for $1$, red for 1-1). We verified that $\omega$ was not a coboundary, confirming that one cannot consistently choose an orientation for the data (since $H^{1}(\mathbb{RP}^{2};\mathbb{Z}_{2})\cong\mathbb{Z}_{2}$, verifying non-triviality is sufficient to uniquely determine the orientation class). With respect to $\omega$, the twisted Euler number of $\Omega$ was found to be $\pm 3$, in agreement with our theoretical model. This confirmed the global topology of the underlying 3-manifold up to homeomorphism.
Using the algorithm described in Section 5, we computed the cobirth and codeath of each characteristic class representative (see Table 3). Note that each representative was found to be a cocycle on the entire nerve of $\mathcal{U}$. The twisted Euler class representative became a coboundary after a single edge was removed from $\mathcal{N}(\mathcal{U})$. This makes sense, because introducing a cut in $\mathbb{RP}^{2}$ yields a space which is homotopy equivalent to $\mathbb{S}^{1}$, hence has no homology in dimension $2$. On the other hand, the orientation class representative only restricted to a coboundary after 110 of the 177 1-simplices in $\mathcal{N}(\mathcal{U})$ were removed. One can interpret this result as a reflection of the compatibly non-orientable structures of the bundle and the base space. More specifically, the restriction of the underlying bundle $(\mathcal{O}_{\text{prism}},\pi,\mathbb{RP}^{2})$ to any non-nullhomotopic loop (e.g. an equator $\mathbb{RP}^{1}\subset\mathbb{RP}^{2}$) is non-orientable, hence isomorphic to the Klein bottle. Thus, the restriction of an orientation class representative $\omega\in Z^{1}(\mathcal{N}(\mathcal{U});\mathbb{Z}_{2})$ to any subcomplex of $\mathcal{N}(\mathcal{U})$ which includes a non-nullhomotopic loop must still be non-trivial.
@@
To illustrate further, we isolated the set of densities $\widetilde{X}_{\text{prism}}\subset X_{\text{prism}}$ whose base point projections satisfied $|z|<0.1$, then projected these base points to the equator $\mathbb{RP}^{1}\subset\mathbb{RP}^{2}$. Using the same cover of $\mathbb{RP}^{1}$ as we used to analyze the optical flow dataset, we recomputed local coordinates and approximate transition matrices and verified that the resulting discrete approximate cocycle produced a non-trivial orientation class representative. After dropping the heaviest edge from the nerve, we used the coordinatization algorithm described in Section 6 to compute global coordinates for $\widetilde{X}_{\text{prism}}$ compatible with the resulting subcomplex.
Figure 18 shows a sample of $\widetilde{X}_{\text{prism}}$ arranged according to assigned base angles in $\mathbb{RP}^{1}$ and assigned fiber angles in $\mathbb{S}^{1}$. Observe that the base angle assigned to each density roughly captures its axis of symmetry, and that this axis steadily traverses a rotation by $\pi$ as the base angle varies from 0 to $\pi$. The fiber angles assigned to densities in any narrow vertical strip roughly parametrize a rotation by $\frac{2\pi}{3}$ about a common axis of symmetry. On the other hand, careful inspection shows that the direction of rotation abruptly flips near the dotted vertical line, reflecting a location in the nerve where a cut was made before coordinatization. This also confirms the Klein bottle structure of the restricted bundle underlying $\widetilde{X}_{\text{prism}}$.
@@
Figure 18: A collection of coordinatized samples from $\widetilde{X}_{\text{prism}}$ arranged according to base projection angle (x) and assigned fiber coordinate (y).
Finally, to demonstrate our bundle map algorithm, we used the procedure described in Section 6 to construct a point cloud embedding $F:X_{\text{prism}}\to V(2,60)\times_{O(2)}\mathbb{S}^{1}\subset\mathbb{R}^{120}$ for the full dataset which respects the global topology implied by the characteristic classes (recall these were well-defined on the entire nerve $\mathcal{N}(\mathcal{U})$). One could potentially reduce the ambient dimension further by applying Principal Stiefel Coordinates to the dataset of frames in $V(2,60)$ used to produce $F$. The mean projection errors for different projection dimensions are shown in Figure 19 .
@@ A Technical Details
A.1 Approximate Isometries Of The Circle
The main result in this section is Lemma 3.10, which roughly states that discrete approximate local trivialization functions on a common domain are approximately related by an isometry of $\mathbb{S}^{1}$. We first review some definitions and introduce technical results which will be used in the proof.
Definition A.1.
Let $d_{\mathbb{S}^{1}}$ denote geodesic distance on $\mathbb{S}^{1}$, and let $d_{\infty}$ be the sup metric on the space $\text{Fun}(\mathbb{S}^{1},\mathbb{S}^{1})$. Note that for $t\in[0,2\pi]$ one has
$d_{\mathbb{S}^{1}}(e^{is},e^{it})=\min\{|s-t|,\,2\pi-|s-t|\}\in[0,\pi]$
The next three lemmas are used in the proof of Proposition A.5:
Lemma A.2.
Suppose $f,g\in\text{Homeo}(\mathbb{S}^{1})$ such that $\deg(f)\neq\deg(g)$. Then, there exists some $z\in\mathbb{S}^{1}$ such that f(z)=g(z)f(z)=-g(z).
Proof.
Let $h=(-g)^{-1}\circ f\in\text{Homeo}(\mathbb{S}^{1})$. Note that multiplication by 1-1 is a degree-1 homeomorphism, so the properties of degree imply
$\deg(h)\quad=\quad\deg((-g)^{-1})\cdot\deg(f)\quad=\quad\deg(-g)\cdot\deg(f)\quad=\quad\deg(g)\cdot\deg(f)\quad=\quad-1$
One can show that every orientation-reversing homeomorphism of $\mathbb{S}^{1}$ has a fixed point (this is an easy consequence of the Intermediate Value Theorem applied to a continuous lift), and if $z_{0}$ is a fixed point of $h$, then $(-g^{-1})\circ f(z_{0})=z_{0}$ (or equivalently, $f(z_{0})=-g(z_{0})$).
Lemma A.3.
Let $(X,d)$ be a metric space. For any fixed $x_{0}\in X$, the assignment $x\longmapsto d(x_{0},x)$ is $1$-Lipschitz:
$|d(x_{0},x)-d(x_{0},x^{\prime})|\leq d(x,x^{\prime})\hskip 28.45274ptx,x^{\prime}\in X$
Proof.
Define $f(x)=d(x_{0},x)$ for all $x\in X$. For any $x,x^{\prime}\in X$, the triangle inequality implies
$d(x,x^{\prime})\geq d(x_{0},x)-d(x_{0},x^{\prime})\hskip 28.45274ptd(x,x^{\prime})\geq d(x_{0},x^{\prime})-d(x_{0},x)$
The result is immediate.
Lemma A.4.
Let $a,b\in\mathbb{S}^{1}$ such that $d_{\mathbb{S}^{1}}(a,b)=\frac{\pi}{2}$. Suppose $z,w\in\mathbb{S}^{1}$ such that
$\max\{\left|d_{\mathbb{S}^{1}}(z,a)-d_{\mathbb{S}^{1}}(w,a)\right|,\quad\left|d_{\mathbb{S}^{1}}(z,b)-d_{\mathbb{S}^{1}}(w,b)\right|\}\quad\leq\quad\delta$
for some $\delta<\frac{\pi}{2}$. Then $d_{\mathbb{S}^{1}}(z,w)\leq\delta$.
Proof.
Since $d_{\mathbb{S}^{1}}$ is $O(2)$-invariant, it suffices to prove the lemma for the case $a=1$ and $b=i$.
Let $s,t\in[0,2\pi)$ and $z=e^{it},w=e^{is}\in\mathbb{S}^{1}$. We first show $d_{\mathbb{S}^{1}}(e^{it},e^{is})\leq\delta$ in the case when $t<\frac{\pi}{2}$ and $s<\pi$. We then argue that the hypothesis $\delta<\frac{\pi}{2}$ implies $e^{is}$ and $e^{it}$ cannot lie in opposite quadrants and use $O(2)$-invariance to conclude that the identity holds for all possible $s,t$.
Suppose $t\in[0,\frac{\pi}{2})$. If $s\in[0,\pi)$, then
$d_{\mathbb{S}^{1}}(e^{it},e^{is})\quad=\quad|t-s|\quad=\quad\left|d_{\mathbb{S}^{1}}(e^{it},1)-d_{\mathbb{S}^{1}}(e^{is},1)\right|\quad\leq\quad\delta$
The case where $s\in[\frac{3\pi}{2},2\pi)$ reduces to the case above by applying a reflection, so $d_{\mathbb{S}^{1}}(e^{it},e^{is})\leq\delta$ in this case by $O(2)$-invariance of $d_{\mathbb{S}^{1}}$. The case where $s\in[\pi,\frac{3\pi}{2})$ is forbidden by the hypothesis $\delta<\frac{\pi}{2}$. To see this, suppose $s\in[\pi,\frac{3\pi}{2})$. Then,
$\delta\quad\geq\quad\left|d_{\mathbb{S}^{1}}(e^{it},1)-d_{\mathbb{S}^{1}}(e^{is},1)\right|\quad=\quad|t-(2\pi-s)|\quad=\quad|2\pi-(s+t)|$ (37)
$\delta\quad\geq\quad\left|d_{\mathbb{S}^{1}}(e^{it},i)-d_{\mathbb{S}^{1}}(e^{is},i)\right|\quad=\quad\left|\left(\frac{\pi}{2}-t\right)-\left(s-\frac{\pi}{2}\right)\right|\quad=\quad|\pi-(s+t)|$ (38)
Since $\delta<\frac{\pi}{2}$, inequalities (37) and (38) are contradictory. We conclude $d_{\mathbb{S}^{1}}(e^{it},e^{is})$ for all feasible $s$ when $t\in[0,\frac{\pi}{2})$. Thus, by $O(2)$-invariance of $d_{\mathbb{S}^{1}}$, the result holds for all possible $s,t\in[0,2\pi)$.